### v1

# Leveraging AI to Accelerate Clinical Data Cleaning: A Comparative Study of AI-Assisted vs. Traditional Methods

Abstract.
Clinical trial data cleaning represents a critical bottleneck in drug development, with manual review processes struggling to manage exponentially increasing data volumes and complexity. This paper presents Octozi, an artificial intelligence-assisted platform that combines large language models with domain-specific heuristics to transform clinical data review. In a controlled experimental study with experienced clinical reviewers (n=10), we demonstrate that AI assistance increased data cleaning throughput by 6.03-fold while simultaneously decreasing cleaning errors from 54.67% to 8.48% (a 6.44-fold improvement). Crucially, the system reduced false positive queries by 15.48-fold, minimizing unnecessary site burden. These improvements were consistent across reviewers regardless of experience level, suggesting broad applicability. Our findings indicate that AI-assisted approaches can address fundamental inefficiencies in clinical trial operations, potentially accelerating drug development timelines and reducing costs while maintaining regulatory compliance. This work establishes a framework for integrating AI into safety-critical clinical workflows and demonstrates the transformative potential of human-AI collaboration in pharmaceutical clinical trials.
copyright: acmlicensedjournalyear: 2025
1. Introduction
The pharmaceutical industry faces an unprecedented data management crisis. Clinical trials now generate three times more data than a decade ago (Taves, 2010), yet the fundamental processes for ensuring data quality remain largely unchanged since the advent of electronic data capture (EDC). This disconnect between data volume and processing capability creates substantial delays in drug development, with each day of database lock delay potentially costing \$500,000 in lost revenue across a product’s lifecycle (Smith et al., 2024). More critically, inadequate data cleaning risks missing crucial safety signals that could impact patient welfare.
Clinical data cleaning, the systematic process of reviewing, validating, and reconciling data collected from investigative sites, serves as the foundation for regulatory submissions to authorities including the US Food and Drug Administration (FDA) and European Medicines Agency (EMA). This process extends beyond simple error detection to encompass complex medical judgment, requiring specialized expertise to identify clinically meaningful discrepancies while maintaining regulatory compliance. The stakes are substantial: data quality issues represent one of the most common causes of regulatory findings, potentially delaying or preventing drug approval.
Traditional approaches to clinical data cleaning rely heavily on manual processes that have remained fundamentally unchanged for decades. Clinical data managers deploy programmatic edit checks to identify basic inconsistencies, while medical monitors conduct labor-intensive reviews of individual patient narratives, laboratory values, and case report forms (CRFs). This manual approach suffers from several critical limitations: inconsistent application of clinical judgment across reviewers, inability to scale with increasing data volumes, and susceptibility to human error during repetitive tasks. Furthermore, the cognitive burden of synthesizing information across multiple data sources often leads to missed relationships and delayed identification of safety signals.
The advent of artificial intelligence, particularly large language models (LLMs) trained on biomedical literature, offers a paradigm shift in clinical data management. These models demonstrate remarkable capability in understanding medical terminology, identifying semantic relationships, and applying consistent logic across vast datasets (Park et al., [n. d.]; Tu et al., 2024). When combined with domain-specific algorithms and clinical heuristics, AI systems can potentially automate routine aspects of data review while augmenting human expertise for complex medical decisions.
Despite this promise, the integration of AI into clinical workflows faces significant challenges. Regulatory bodies require transparent, auditable processes for safety-critical decisions. Clinical reviewers express concerns about ”black-box” algorithms making medical judgments. Perhaps most importantly, any AI system must demonstrate not only efficiency gains, but also maintenance or improvement of data quality standards to gain acceptance in this highly regulated environment. Here, we present Octozi, an AI-assisted clinical data cleaning platform designed to address these challenges through a hybrid approach combining machine learning with expert-encoded clinical logic. Our system leverages unified data pipelines to continuously ingest and contextualize diverse clinical trial data, presenting AI-generated insights through an intuitive interface that maintains human oversight while dramatically improving efficiency. We hypothesized that this human-AI collaboration model would improve both the speed and accuracy of clinical data review compared to traditional methods.
To rigorously test this hypothesis, we conducted a controlled experimental study with experienced clinical reviewers, comparing the performance between traditional spreadsheet-based review and AI-assisted workflows. Our results demonstrate that AI assistance fundamentally can shift the focus of clinical development teams from reactive data cleaning to proactive safety and efficacy monitoring. By dramatically reducing the time spent on routine data reconciliation, AI enables teams to engage in higher value activities: identifying safety signals earlier, optimizing trial protocols in real time, and making more informed decisions about patient care. This transformation not only accelerates database-lock but enables more adaptive, responsive trials that can pivot based on emerging data patterns, ultimately reducing both development risk and time to market for life-saving therapies.
2. Methods
2.1. Study Design and Statistical Considerations
We designed a within-subjects controlled experiment to evaluate the impact of AI assistance on clinical data review performance. This design allowed each participant to serve as their own control, minimizing inter-individual variability while maximizing statistical power. Prior to the study, we conducted an a priori power analysis to determine an appropriate sample size for detecting a meaningful improvement in clinical data cleaning throughput. Assuming a conservative large effect size (Cohen’s d = 1.0), a two-sided paired t-test with $\alpha$ = 0.05 would require a minimum of 10 participants to achieve 80% statistical power. This threshold was selected to detect at least a 50% improvement in throughput, which we considered a clinically meaningful gain based on early pilot observations.
Upon completion of the study, the observed effect size was substantially larger than anticipated. Based on the actual performance data, the mean improvement in throughput between baseline and AI-augmented methods was 17.1 records per session (standard deviation = 8.16), yielding an observed Cohen’s d of 2.10. A post hoc power analysis using these values confirmed that our sample of 10 participants yielded a power greater than 99.9% to detect a statistically significant difference in performance ($\alpha$ = 0.05, two-sided). This result confirms that the study was more than adequately powered to evaluate the efficacy of the AI-assisted system.
Study participants were recruited from a diverse pool of clinical research professionals through targeted outreach to pharmaceutical companies, contract research organizations, and academic medical centers. Eligibility criteria included: (1) minimum two years of experience in clinical data review roles, (2) proficiency in adverse event adjudication, and (3) familiarity with oncology clinical data cleaning. Pre-study assessments captured detailed professional backgrounds including clinical data cleaning experience, familiarity with specific data management systems, and work experience.
Table 1. Summary of clinically meaningful discrepancy categories introduced into the synthetic dataset. These six categories were identified through expert consultation as the most common adverse event-related data quality issues encountered during oncology clinical trial reviews.
Discrepancy categories
1 Inappropriate concomitant medication to treat an adverse event
2 Timing of concomitant medication administration and adverse event do not align
3 Incorrect severity score attached to an adverse event based on the description of the event
4 Mismatched dosing change
5 Incorrect causality assessment of adverse event
6 No supporting data for adverse event
2.2. Synthetic Dataset Construction and Validation
The experimental datasets were derived from a comprehensive phase III oncology trial database containing electronic data capture (EDC) information across 51 separate case report form datasets for over 150 patients. Given our focus on adverse event discrepancy detection, we strategically selected a subset of 8 CRFs that were directly relevant to adverse event assessment and documentation. This focused approach ensured that participants could concentrate on clinically meaningful data relationships without being overwhelmed by extraneous information.
To create synthetic patients while preserving complete anonymity, we utilized a library-based refinement generation approach. We first constructed comprehensive libraries of clinical elements by extracting all unique adverse events, concomitant medications, ancillary procedures, and medical history entries from the original dataset. These libraries preserved the diversity and complexity of real clinical data while completely dissociating individual data points from their original patient contexts. For each synthetic patient, we randomly sampled entries from these libraries, with sampling frequencies weighted by their occurrence rates in the original data to maintain realistic clinical patterns.
The initial synthetic data underwent rigorous clinical refinement by expert annotators who reviewed each generated patient profile for medical coherence. These experts made targeted modifications to enhance realism, ensuring that temporal relationships, severity progressions, and treatment patterns reflected authentic clinical scenarios. Laboratory values and vital signs were particularly carefully constructed, while initially set to normal ranges, these values were systematically modified to align with documented adverse events. For instance, patients with documented anemia had their hemoglobin values adjusted to clinically appropriate low levels, maintaining internal consistency throughout the dataset.
To simulate realistic data cleaning scenarios, we systematically introduced clinically meaningful discrepancies based on extensive consultation with industry experts. Through structured interviews with three senior clinical scientists and medical monitors, we identified the most common and impactful data quality issues encountered in oncology trials. These interviews revealed six primary categories of discrepancies that represent the majority of clinically significant data quality issues in modern trials (detailed in Table 1). Discrepancies were introduced into 10% of all data points using a stratified randomization approach, ensuring equal distribution across both experimental conditions and discrepancy types.
2.3. Experimental Protocol
The study protocol was designed to minimize bias and consisted of three carefully structured phases. The baseline assessment phase began with participants receiving standardized training on the dataset structure and common discrepancy types. Following a 10-minute briefing period, participants conducted manual reviews using industry-standard spreadsheet tools, with time tracking initiated immediately after the orientation and participant setup of their spreadsheet tool environment. This approach replicated typical clinical data management workflows used across the pharmaceutical industry.
The AI-assisted review phase commenced with an interactive tutorial on the Octozi platform features, including live demonstrations of the system’s capabilities and hands-on practice with a specially designed training dataset containing known discrepancies. After confirming proficiency with the platform, participants were allotted 30 minutes to review a matched complexity dataset using Octozi’s full functionality.
The final phase consisted of standardized assessment using two validated instruments. The NASA Task Load Index (NASA-TLX) was administered to quantify the cognitive workload associated with each review method across six dimensions: mental demand, physical demand, temporal demand, performance, effort, and frustration. This multidimensional assessment was particularly relevant given the cognitive intensity of clinical data review and the potential for AI to reduce mental burden. The System Usability Scale (SUS) provided a standardized measure of the Octozi platform’s usability, generating a composite score that enabled benchmarking against industry standards for enterprise software. Together, these instruments captured both the cognitive impact of AI assistance and the practical usability of the technology, providing crucial insights into the feasibility of widespread adoption in clinical research settings.
Figure 1. Individual participant throughput in correctly cleaned clinical data listings using traditional spreadsheets versus the AI-assisted Octozi platform. Across all participants, Octozi substantially increased data cleaning throughput, with a mean improvement of 6.03-fold (range: 3.7x to 15x). Participants cleaned a mean of 3.4 data points (range: 1–6) with spreadsheets and a median of 20.5 data points (range: 7–35) using Octozi within the same 30-minute review period. Improvements were observed consistently across all participants, highlighting the effectiveness of AI assistance in clinical data review tasks.
2.4. Discrepancy Detection Algorithms
Octozi employs proprietary large language models (LLMs) derived from Llama 4, specifically fine-tuned for clinical trial data cleaning tasks. The discrepancy detection engine combines these LLMs with a suite of heuristic algorithms, systematically developed in collaboration with expert clinical data reviewers. The hybrid approach enables the platform to efficiently identify complex data inconsistencies by integrating domain-specific clinical reasoning with advanced language model capabilities. When evaluated on the annotated synthetic dataset, the discrepancy detection algorithms achieved a classification accuracy of 83.6%, with a recall of 97.5% and precision of 77.2%. Importantly, the accuracy metric reflects correct prediction of the specific discrepancy category, rather than simple anomaly detection, underscoring the model’s capacity to replicate nuanced clinical decision-making patterns.
3. Results
3.1. Data Cleaning Efficiency
Analysis of individual participant performance revealed consistent and substantial improvements across all reviewers (Fig. 1). Using traditional spreadsheet methods, participants correctly cleaned a mean of 3.4 data points (range: 1-6), while with Octozi assistance, this increased to a median of 20.5 data points (range: 7-35) within the same 30-minute review period. The improvement ratios ranged from 3.7x to 15x, with a mean improvement of 6.03x across all participants.
Notably, the visualization of participant trajectories (Fig. 2) demonstrates that every single participant experienced both improved accuracy and increased throughput when using Octozi, with no instances of speed-accuracy tradeoffs. The consistent upward and rightward movement of all participant data points illustrates the dual benefit of AI assistance, faster and more accurate data cleaning. The worst performer using Octozi performed better than the best participant using spreadsheets, highlighting the substantial and consistent improvement.
Figure 2. Accuracy versus throughput for individual participants using traditional spreadsheets and the Octozi AI-assisted platform. Each arrow represents the change in performance for a single participant, demonstrating consistent improvements in both accuracy and throughput with Octozi. On average, participants achieved 92% accuracy while cleaning six times more data compared to 46% accuracy with spreadsheets. Notably, no participant experienced a trade-off between speed and accuracy; all participants improved on both dimensions, with the lowest Octozi performance exceeding the highest spreadsheet performance.
3.2. Doubling of Correct Discrepancy Accuracy
The confusion matrices reveal substantial improvements to clinical data cleaning in terms of classification accuracy (Fig. 3). Overall accuracy improved from 45.3% with manual review to 91.5% with Octozi assistance, representing a 46.2 percentage point absolute improvement or a 2.02-fold relative improvement. This enhancement was driven by dramatic improvements in precision, which increased from 44.6% to 93.2% (48.6 percentage point improvement), indicating that AI assistance nearly eliminated false positive identifications.
The F1-score, which balances precision and recall, improved from 58.5 to 91.0, demonstrating that AI assistance enhanced both the completeness and accuracy of discrepancy detection. Interestingly, recall remained high in both conditions (85.2% vs 88.9%), suggesting that experienced reviewers were already adept at recognizing potential issues, but AI assistance helped them correctly distinguish true discrepancies from false alarms.
3.3. Reduction in False Positive Queries: Minimizing Site Burden
Perhaps most significantly for operational efficiency, the confusion matrices demonstrate a dramatic reduction in false positive classifications. With spreadsheet-based review, 36 out of 75 total classifications (48.0%) were false positives—instances where reviewers incorrectly flagged clean data as discrepant. With Octozi, this dropped to just 7 out of 224 classifications (3.1%), representing a 15.48-fold reduction in false query generation.
This reduction has profound implications for trial operations. In the manual condition, nearly half of all queries sent to sites would have been unnecessary, creating substantial bottlenecks to overburdened sites. The Octozi platforms’s ability to maintain high sensitivity while dramatically improving specificity addresses a fundamental challenge in clinical data management.
3.4. Statistical Significance and Effect Sizes
All primary performance differences between methods achieved statistical significance with large effect sizes. The accuracy improvement yielded a Cohen’s d of 2.8, indicating an extremely large effect size that represents meaningful practical significance beyond statistical significance. Similarly, throughput improvements demonstrated effect sizes exceeding 3.0, confirming that observed differences reflect genuine performance advantages rather than statistical artifacts. Paired t-test analysis confirmed that improvements were consistent across participants (p ¡ 0.001 for all primary measures), with 95% confidence intervals that excluded the possibility of equivalent performance between methods. The magnitude and consistency of observed improvements provide strong evidence for the superiority of AI-assisted clinical data review in both efficiency and accuracy domains.
3.5. Usability and User Experience
Post-study usability assessments revealed unanimous preference for the AI-assisted system across all evaluated dimensions. NASA-TLX scores indicated significantly reduced perceived workload with the automated approach, with participants reporting lower mental demand, temporal pressure, effort requirements, and frustration levels compared to traditional methods. System Usability Scale scores for the AI system were on average 86.3, indicating excellent usability, while traditional spreadsheet methods scored 56 on average, reflecting substantial user experience challenges.
Qualitative feedback consistently highlighted the AI system’s ability to prioritize findings, provide contextual clinical information, and reduce the cognitive burden associated with cross-referencing multiple data sources. Participants particularly valued the system’s capability to surface complex cross-domain discrepancies that would be difficult to identify through manual review. Several participants noted that the AI system helped them identify discrepancy types they had not previously recognized, suggesting potential educational benefits beyond immediate productivity gains.
Figure 3. Confusion matrices and performance metrics comparing manual spreadsheet review and AI-assisted review using Octozi. Octozi substantially improved classification performance, increasing overall accuracy from 45.3% to 91.5% (+46.2 percentage points) and precision from 44.6% to 93.2% (+48.6 percentage points), while maintaining high recall. The F1-score improved from 58.5 to 91.0, reflecting balanced gains in precision and recall. Notably, false positive classifications—clean data incorrectly flagged as discrepancies—were reduced 15-fold (48.0% to 3.1%), indicating a significant reduction in erroneous queries and potential site burden.
4. Discussion
Our findings demonstrate that AI assistance can fundamentally transform clinical trial data cleaning from a rate-limiting bottleneck into an accelerator of drug development. The magnitude of improvement observed, over 6-fold increase in throughput coupled with doubling of accuracy, exceeds typical incremental advances in clinical trial operations and suggests a true paradigm shift in how data quality assurance can be conducted.
4.1. Transforming Clinical Trial Timelines Through Enhanced Data Review Efficiency
The demonstrated 6-fold improvement in data review throughput achieved through AI-assisted clinical data cleaning has profound implications for accelerating clinical trial timelines across multiple critical phases. Traditional data cleaning bottlenecks that routinely delay trial milestones by weeks or months could be substantially reduced. This acceleration is particularly crucial for rare disease trials where patient populations face limited therapeutic options and urgent medical needs that make time-to-market considerations a matter of life and death.
The efficiency gains observed in this study directly address the most time-sensitive aspects of clinical trial conduct. Database preparation for interim analyses, which currently requires 4-8 weeks of intensive data cleaning effort, could potentially be completed within 1-2 weeks using AI-assisted methods. This compression of data review timelines enables more frequent and timely interim analyses, allowing Data Safety Monitoring Boards to make critical go/no-go decisions with more current information and potentially identifying safety signals or efficacy trends earlier in the trial lifecycle. For adaptive trial designs that rely on interim data to modify study parameters, the ability to rapidly prepare clean datasets becomes essential for maintaining trial momentum and maximizing the value of adaptive design features.
4.2. Revolutionizing Database Lock Procedures and Regulatory Submission Timelines
Database lock represents one of the most critical bottlenecks in running a clinical trial, where comprehensive data cleaning must achieve near-perfect quality standards before final analysis and regulatory submission. The 46.2% improvement in accuracy demonstrated by AI-assisted review could reduce the iterative cycles of query generation and resolution that traditionally extend database lock timelines by weeks. Current industry practice typically requires 7 weeks between last patient visit and database lock (Harper et al., 2021). AI-assisted review could potentially reduce this timeline to 1-2 weeks, representing a 75% reduction in database lock duration.
This acceleration has cascading effects throughout the regulatory submission process. Earlier database lock enables earlier statistical analysis, regulatory document preparation, and submission filing, potentially advancing market approval timelines by 3-6 months. For breakthrough therapies addressing unmet medical needs, this timeline compression could translate to thousands of additional patients gaining access to life-saving treatments months earlier than would be possible with traditional data management approaches. The economic implications are equally significant, with reduced development timelines extending patent life and improving return on investment for pharmaceutical sponsors.
4.3. Enabling Dynamic and Adaptive Trial Designs
The efficiency and accuracy improvements demonstrated by AI-assisted data review create new possibilities for implementing dynamic trial designs that were previously impractical due to data management constraints. Real-time data monitoring and continuous data cleaning become feasible when AI systems can maintain high-quality datasets with minimal human intervention. This capability enables seamless implementation of adaptive randomization, dose escalation decisions, and futility assessments that require rapid access to clean, analysis-ready data.
Platform trials and master protocols, which evaluate multiple treatments or indications simultaneously, could particularly benefit from AI-assisted data review. These complex trial designs generate massive datasets with intricate cross-study dependencies that challenge traditional manual review methods. The AI system’s demonstrated ability to identify cross-domain discrepancies and maintain consistency across multiple data sources makes it ideally suited for managing the complex data relationships inherent in platform trial designs. This could accelerate the adoption of more efficient trial designs that evaluate multiple therapeutic approaches in parallel rather than sequential monotherapy studies.
4.4. Reallocating Human Expertise to Higher-Value Clinical Activities
The dramatic efficiency improvements achieved through AI-assisted data review create opportunities to reallocate scarce clinical expertise from routine data checking activities to higher-value clinical interpretation and decision-making tasks. Clinical data managers and medical monitors can redirect their time from manual discrepancy detection to clinical assessment of identified issues, protocol deviation investigation, and strategic data quality planning. This shift from mechanical data review to clinical reasoning maximizes the value of expensive clinical expertise while maintaining rigorous oversight of automated processes.
The educational benefits observed in this study, where participants identified previously unrecognized discrepancy types, suggest that AI systems can enhance rather than replace human clinical judgment. By systematically flagging complex patterns and providing contextual clinical information, AI-assisted review can help clinical professionals develop deeper expertise in data quality assessment while simultaneously improving review efficiency. This human-AI collaboration model ensures that the clinical knowledge and regulatory experience essential for trial success remain central to data management processes while leveraging automation to eliminate routine, time-consuming tasks.
5. Conclusion
The integration of AI into clinical trial data cleaning represents a transformative opportunity to address fundamental inefficiencies in drug development. Our controlled experimental study provides robust evidence that AI assistance can simultaneously improve the speed, accuracy, and consistency of clinical data review while reducing operational burden on clinical trial sites.
These findings have immediate practical implications for the pharmaceutical industry, suggesting that AI adoption could accelerate drug development timelines by weeks to months while improving data quality and reducing costs. More broadly, our results demonstrate the potential for human-AI collaboration to enhance performance in complex, safety-critical domains requiring both efficiency and expert judgment.
As the pharmaceutical industry grapples with increasing trial complexity and data volumes, AI-assisted approaches offer a scalable solution that enhances rather than replaces human expertise. The consistent benefits observed across reviewer experience levels suggest that AI can democratize access to high-quality data review capabilities, potentially enabling more efficient global clinical trials.